OpenStudy (anonymous):

@kelliegirl33 plz plz help i will fan and medal

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

what ya got ?

OpenStudy (anonymous):

Corrine plotted 3 points from the equation x = 4 on this coordinate grid. She drew a straight line through the points. Which ordered pair would also be on this line? A. (+6, +2) B. (+6, +4) C. (+4, +2) D. (+2, +4)

OpenStudy (anonymous):

there is the graph

OpenStudy (anonymous):

look at your answer choices....which answer also has x = 4 because , being a vertical line, all the answers that are on the line will have x being 4

OpenStudy (anonymous):

so um uh

OpenStudy (anonymous):

its C.....because answer choice C has x being 4

OpenStudy (anonymous):

Brian plotted 3 points from the equation y = 4 on this coordinate grid. He drew a straight line through the points. Which ordered pair would also be on this line? A. (+4, +1) B. (+7, +4) C. (+5, +2) D. (+5, +5)

OpenStudy (anonymous):

which answer choice has y also being 4 ?

OpenStudy (anonymous):

que

OpenStudy (anonymous):

Its B

OpenStudy (anonymous):

Steve plotted 3 points from the equation y = 3 × x on this coordinate grid. He drew a straight line through the points. Which ordered pair would also be on this line? A. (+5, +15) B. (+4, +7) C. (+5, +3) D. (+12, +4)

OpenStudy (anonymous):

?hello?

OpenStudy (anonymous):

I think its gonna be A

OpenStudy (anonymous):

you sure

OpenStudy (anonymous):

i mean my gma said i need to hurry with my work cuz its gonna be lighting and thundering and a flash flood so

OpenStudy (anonymous):

I think so...because the slope is 3 and if it keeps going up 3 and over 1. it will hit (5,15)

OpenStudy (anonymous):

Jaime plotted 3 points from the equation y = x + 5 on this coordinate grid. He drew a straight line through the points. Which ordered pair would also be on this line? A. (+7, +2) B. (+2, +7) C. (+8, +3) D. (+3, +5)

OpenStudy (anonymous):

sub in your answer choices... y = x + 5 choice B (2,7) ..x = 2 and y = 7 y = x + 5 7 = 2 + 5 7 = 7 (correct) answer is B

OpenStudy (anonymous):

Ray plotted 3 points from the equation y = 3 × x + 2 on this coordinate grid. He drew a straight line through the points. Which ordered pair would also be on this line? A. (+3, +11) B. (+3, +15) C. (+4, +10) D. (+4, +13)

OpenStudy (anonymous):

y = 3x + 2 choice A ... (3,11) x = 3 and y = 11 11 = 3(3) + 2 11 = 9 + 2 11 = 11 (correct)

OpenStudy (anonymous):

its

OpenStudy (anonymous):

you do not even need the graphs...just sub in your answer choices and see which one satisfies the equation

OpenStudy (anonymous):

last one is A

OpenStudy (anonymous):

100%

OpenStudy (anonymous):

cool :)